The average speed of one airplane is 200 mph faster than that of an automobile. If the automobile starts at 3 am and the airplane starts from the same place at 11 am, they will have traveled the same number of miles by 1 pm. Find the rate of the automobile.
___mph

Answer by PBMathandscience(9)

Let the speed of the car be "X". From 3 am until 1 pm (10 hours), the car will travel 10*X miles.
The speed of the airplane is (X +200). From 11 am until 1 pm (2 hours), the plane will travel 2*(X+200).
The distance travelled is the same, so:
10*X = 2*(x+200)= 2x+400
Subtract 2*X from both sides:
8*X = 400
X = 400/8 = 50 mph.